OpenStudy (anonymous):

whats the surface area of a cylinder with a diameter of 40 and a height of 60. Use 3.14 for π. we have to Round our answer to the nearest tenth.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

SA=2(pi)(r^2) + 2(pi)(r)(h) r=1/2d r=20 2*3.14*20*20+2*3.14*20*60 10048.0